OpenStudy (anonymous):

The expression 7! is equal to which of the following? A. 7 B. 720 C. 5040 D. 40,320

OpenStudy (anonymous):

kwon

OpenStudy (anonymous):

7 is equal to 7

OpenStudy (anonymous):

Did I read the problem wrong?

OpenStudy (anonymous):

ummm im not sure im not good with math so

OpenStudy (anonymous):

the problem is 7 is equal to what? what is that thing after the 7?

OpenStudy (anonymous):

the following

OpenStudy (anonymous):

I am pretty sure that it is A. I wish i was taking this kind of math.

OpenStudy (anonymous):

not me i hate it...ok thanks

OpenStudy (johnweldon1993):

7! means 7 factorial...what you do with that "!" sign...is you multiply 7 by every number before it...i.e \[\large 7\times 6 \times 5 \times 4 \times 3 \times 2 \times 1 = ? \]

OpenStudy (anonymous):

I am glad you know that because I have never seen the !

OpenStudy (anonymous):

i dont know why thats there

OpenStudy (johnweldon1993):

I believe I learned that in Algebra 1? Long time ago so not too sure when but definitely know it :)

OpenStudy (johnweldon1993):

Why what is there? @brooke112heart ? That '!' sign IS used in mathematics ...and it is a factorial as stated above :)

OpenStudy (anonymous):

um to be onest idk.... ok thanks you guys....bye

OpenStudy (anonymous):

good job @johnweldon1993

OpenStudy (anonymous):

bye